# Leveraging Semantic Attribute Binding for Free-Lunch Color Control in Diffusion Models

### v1

@@ 4.2 Qualitative Comparisons
To demonstrate the versatility and effectiveness of ColorWave , we present qualitative results across several key capabilities: precise color attribution across diverse objects, fine-grained color control, and generalization to complex color patterns and textures. These experiments highlight the flexibility of our training-free approach in scenarios that would require extensive additional training in previous methods. More results can be found in the Supplementary.
Arbitrary color attribution. We evaluate ColorWave ’s ability to precisely apply user-specified colors to a wide range of subjects including inanimate objects, animals, plants, and human clothing. Figure 6 demonstrates that our method successfully generates images where target objects accurately reflect the desired reference colors. Notably, ColorWave maintains high visual quality across diverse scenarios, preserving natural lighting effects, surface properties, and contextual integration while achieving precise color matching.
@@ 4.3 Quantitative Analysis
We compare ColorWave with training-based and training-free baselines. Such quantitative results are demonstrated in Table 1. For each method, we generated images, and extract the mask of the object, following the same pipeline as in ColorPeel [4]. Percentages in MAE metrics denote the percentage of pixels inside the mask used for the computation (selecting those closest to the ground truth). This table clearly shows the superiority of ColorWave . Particularly, ColorWave achieved notably lower $\Delta E$ error in CIE Lab color space as compared to the existing training-free methods and it is only worse than the training-based state-of-the-art ColorPeel, which indicates that ColorWave generates perceptually better colors than its direct competitors. In addition, ColorWave also achieved comparatively much lower mean angular error in both sRGB and Hue when compared to the other training-free methods, which signifies a higher degree of color accuracy in terms of chromaticity and hue in our generated images.
Table 1: Quantitative comparison with baselines over various evaluation metrics. All numbers are the smaller the better ($\downarrow$). The best results of both training-based and training-free technique streams are highlighted in bold. Training time is provided in the last column.
@@
Training Based TI [10] 48.98 44.29 15.22 19.51 23.90 52.66 69.35 90.88 118
DB [38] 50.71 46.29 14.75 19.30 23.70 47.12 67.13 88.72 56
CD [21] 48.47 42.23 13.43 17.93 22.43 31.63 55.07 78.43 24
ColorPeel (3D) [4] 21.39 16.51 4.36 7.76 12.08 2.63 6.47 21.35 19
ColorPeel (2D) [4] 20.45 15.29 4.83 7.88 12.13 3.18 7.43 21.46 19
Training Free SD [35] 47.45 41.55 12.89 20.04 26.93 30.17 54.14 86.38 0
Rich-Text [11] 36.62 32.48 9.91 13.29 18.53 50.55 72.77 93.51 0
ColorWave (Ours) 29.91 26.57 8.65 10.39 12.71 13.69 22.14 33.12 0
User study We conducted a user study with 15 participants to perceptually evaluate our results, comparing ColorWave against ColorPeel [4], TI [10], Rich-Text [11], DB [38], and CD [21]. We followed the same experimental paradigm as the one defined in ColorPeel.
The experiment took place in a controlled lab environment to ensure reliability. All participants were tested for correct color vision using the Ishihara test. The study followed a two-alternative forced choice (2AFC) method. Observers viewed three images on a monitor set to RGB: the central image represented the target color, while the left and right images displayed results generated by our method and one of the competing methods, with their positions randomized. We tested the same 10 different prompts and four colors (red, green, blue, and yellow) defined in ColorPeel.
@@
Figure 8: User study. Z-scores are higher for better human preference. Our method outperforms baselines and the state-of-the-art color control generation method ColorPeel [4].
@@
4.4 Ablations
Color name attribution.
We investigate the constraints of semantic attribute binding by intentionally creating discrepancies between the color mentioned in the prompt and the color specified as input to our model. Figure 9 illustrates these results using a consistent prompt “A photo of a red parrot” while varying the input color values. Our findings reveal a semantic proximity effect, understood by Fig. 3; the success of color attribution depends on the perceptual and semantic distance between the prompt color and the target color. As shown in the upper row of Figure 9, when the target color (orange) shares perceptual similarity with the prompt color (red), our method successfully binds the new color to the parrot. However, when the target color (green) represents an opponent color on the color wheel, the binding fails, and the model defaults to generating a parrot closer to the prompted red color. The lower row demonstrates this effect across a broader range of colors. Pink, which shares perceptual properties with red (both are warm colors with similar hue angles), successfully binds to the parrot. In contrast, semantically distant colors such as olive, navy, and yellow fail to override the prompted “red” descriptor, resulting in parrots that retain reddish tones despite the different color input.
4.5 Limitations
Despite the effectiveness of ColorWave in achieving training-free color control, several limitations remain. First, our approach relies on the quality of cross-attention maps generated by the underlying diffusion model. When these maps exhibit attention leakage—where attention for specific objects spreads to unintended regions—our method’s color attribution precision diminishes. While state-of-the-art diffusion models continue to improve attention mechanism quality, certain complex scenes or ambiguous prompts may still produce suboptimal attribution maps.
Furthermore, our current implementation faces challenges when attributing different colors to multiple objects in a single scene. The cross-attention maps tend to degrade in quality as the number of attribution targets increases, potentially leading to color bleeding between objects. This limitation becomes particularly evident in dense scenes with numerous objects requiring distinct color specifications.
For future works, we intend to address these limitations by exploring robust attention map generation techniques and developing specialized approaches for multi-object color attribution.
@@ 5 Conclusion
We introduced ColorWave , the first training-free approach for precise color control in text-to-image (T2I) diffusion models. By leveraging the latent binding between visual attributes and semantic representations within the IP-Adapter, we enabled direct RGB-level color manipulation without requiring additional model fine-tuning. Our method effectively rewires cross-attention mechanisms to establish precise color attribution while maintaining the generative quality and diversity of pretrained diffusion models. Extensive evaluations demonstrated that our method significantly outperforms prior training-free approaches in color accuracy and object-specific fidelity. Our method also shows competitive performance versus state-of-the-art, fine-tuning methods like ColorPeel, thus overcoming the limitations of prompt-based color descriptions which are computationally expensive. Our findings reveal previously untapped capabilities in existing diffusion models, paving the way for new directions in controllable generation.
Acknowledgements
We acknowledge project PID2022-143257NB-I00, financed by MCIN/AEI/10.13039/501100011033 and FEDER, and the Generalitat de Catalunya CERCA Program. JVC was funded by Grant PID2021-128178OB-I00 funded by MCIN/AEI/10.13039/501100011033, ERDF ”A way of making Europe” and the Departament de Recerca i Universitats from Generalitat de Catalunya with ref. 2021SGR01499. This work was partially supported by the grant Càtedra ENIA UAB-Cruïlla (TSI-100929-2023-2) from the Ministry of Economic Affairs and Digital Transition of Spain.
@@ Appendix A Color binding regions as a function of steps
Figure 10 displays color binding regions based on $\left\lceil\mathbf{A}_{o}\right\rceil$ every 10-th denoising step,
Figure 10: Color binding regions based on $\left\lceil\mathbf{A}_{o}\right\rceil$. Depicted every 10-th denoising step.
Appendix B Experimental details
Evaluation metric details For each generated image, we use the Segment-Anything model [20] to automatically extract masks for the target objects, allowing precise measurement of color attributes only within relevant regions. We report these metrics under multiple conditions: considering all pixels within the object mask (100%), and also considering the top 10% and 50% of pixels closest to the target color, which helps account for natural variations in object appearance such as highlights and shadows.
Implementation details We implement ColorWave using a pretrained Stable Diffusion XL (SDXL) [33] model as our base architecture. For the image conditioning component, we utilize the IP-Adapter framework [51] with an encoder based on OpenCLIP-ViT-H-14 [17]. We only inject the color embedding into the first decoder layer to compute cross-attention maps, which shows better stylization performance as previous works proved [45, 46, 1].
@@ Appendix B Experimental details
Appendix C Extra results of ColorWave
Extra results are showcased in Figures 11 to 18.

### v2

@@ 4.2 Qualitative Comparisons
To demonstrate the versatility and effectiveness of ColorWave, we present qualitative results across several key capabilities: precise color attribution across diverse objects, fine-grained color control, and generalization to complex color patterns and textures. These experiments highlight the flexibility of our training-free approach in scenarios that would require extensive additional training in previous methods. More results can be found in Section F.
Arbitrary color attribution. We evaluate ColorWave ’s ability to precisely apply user-specified colors to a wide range of subjects including inanimate objects, animals, plants, and human clothing. Figure 6 demonstrates that our method successfully generates images where target objects accurately reflect the desired reference colors. Notably, ColorWave maintains high visual quality across diverse scenarios while achieving precise color matching.
@@ 4.3 Quantitative Analysis
We compare ColorWave with training-based and training-free baselines. Such quantitative results are demonstrated in Table 1. For each method, we generated images, and extract the mask of the object, following the same pipeline as in ColorPeel [butt2025colorpeel]. Percentages in MAE metrics denote the percentage of pixels inside the mask used for the computation (selecting those closest to the ground truth). This table clearly shows the superiority of our method. Particularly, ColorWave achieved notably lower $\Delta E$ error in CIE Lab color space as compared to the existing training-free methods and it is only worse than the training-based state-of-the-art ColorPeel, which indicates that ColorWave generates perceptually better colors than its direct competitors. In addition, ColorWave also achieved comparatively much lower mean angular error in both sRGB and Hue when compared to the other training-free methods, which signifies a higher degree of color accuracy in terms of chromaticity and hue in our generated images. While ColorPeel achieves superior numerical performance on some metrics due to its color-specific training, our training-free approach offers significant practical advantages: immediate RGB specification without optimization, broader architectural compatibility, and elimination of per-color training requirements.
Table 1: Quantitative comparison with baselines over various evaluation metrics. All metrics the smaller the better ($\downarrow$). The best results of both training-based and training-free technique streams are highlighted in bold. Training time is provided in the last column.
@@
Training Based TI [textual_inversion] 48.98 44.29 15.22 19.51 23.90 52.66 69.35 90.88 118
DB [ruiz2022dreambooth] 50.71 46.29 14.75 19.30 23.70 47.12 67.13 88.72 56
CD [kumari2022customdiffusion] 48.47 42.23 13.43 17.93 22.43 31.63 55.07 78.43 24
ColorPeel (3D) [butt2025colorpeel] 21.39 16.51 4.36 7.76 12.08 2.63 6.47 21.35 19
ColorPeel (2D) [butt2025colorpeel] 20.45 15.29 4.83 7.88 12.13 3.18 7.43 21.46 19
ColorPeel (3D) [butt2025colorpeel] SD-XL 29.01 23.65 5.46 8.19 12.59 9.98 26.28 51.71 41
ColorPeel (2D) [butt2025colorpeel] SD-XL 28.56 23.25 5.31 8.06 12.45 9.28 25.24 50.53 41
Training Free SD [Rombach_2022_CVPR_stablediffusion] 47.45 41.55 12.89 20.04 26.93 30.17 54.14 86.38 0
Rich-Text [ge2023richtext] 36.62 32.48 9.91 13.29 18.53 50.55 72.77 93.51 0
Rich-Text++ [ge2025expressive] SD-XL 60.31 57.04 21.42 24.35 27.13 81.33 115.08 139.72 0
ColorWave (Ours) SD-XL 22.81 19.64 5.89 7.83 11.00 14.94 23.12 39.68 0
ColorWave (Ours) SD3.5-L 16.04 12.25 5.01 7.78 11.05 3.05 8.44 27.80 0
ColorWave (Ours) FLUX.1-dev 23.04 18.91 5.64 8.64 12.10 10.57 27.96 49.20 0
@@
Figure 8: User study. Z-scores are higher for better human preference. Our method outperforms baselines and the state-of-the-art color control generation method ColorPeel [butt2025colorpeel].
4.4 Ablations
Comparison to IP-Adapter. Direct IP-Adapter application for color control using solid color patches or colored objects suffers from significant limitations that ColorWave addresses. Figure 9 and Table 2 demonstrate naive IP-Adapter usage exhibits substantial color leakage (36.03% of non-target pixels affected vs. 8.80% for ColorWave ), where colors bleed into unintended regions beyond the target object, and produces flat, unrealistic appearances lacking natural lighting variations, shadows, and material properties. The leakage metric definition can be found in Section D. Additionally, direct IP-Adapter generates lower quality images as measured by no-reference metrics NIQE [6353522] (6.17 vs. 4.78) and BRISQUE [6272356] (29.76 vs. 21.49). However, to address the photorealism limitations inherent in direct IP-Adapter application, our method introduces spatial prior addition (Equation 4), which constrains color application through attention map masking, preventing unwanted leakage and preserving natural lighting variations, shadows, and material properties. Despite these improvements in spatial control and visual realism, ColorWave maintains similar color accuracy to the direct approach, demonstrating that our method successfully addresses IP-Adapter’s fundamental limitations for precise color control tasks while enhancing overall generation quality.
Table 2: Comparison between direct IP-Adapter usage and ColorWave for color control tasks. All metrics the smaller the better ($\downarrow$).
Method
Color
leakage
NIQE BRISQUE
IP-Adapter 36.03 6.17 29.76
ColorWave 8.80 4.78 21.49
Figure 9: Comparison between direct IP-Adapter usage and ColorWave. (top) Naive IP-Adapter exhibits flat appearance and significant color leakage. (bottom) ColorWave demonstrates realistic lighting, shadows, and material properties while maintaining precise color control without leakage.
Color name attribution. We investigate semantic attribute binding constraints by creating discrepancies between prompt colors and specified input colors. Figure 10 shows results using consistent prompt “A photo of a red parrot” while varying input colors. Our findings reveal a semantic proximity effect, understood by Fig. 3; the success of color attribution depends on the perceptual and semantic distance between the prompt color and the target color. As shown in Figure 10 top, when the target color (orange) shares perceptual similarity with the prompt color (red), our method successfully binds the new color to the parrot. However, when the target color (green) represents an opponent color on the color wheel, the binding fails, and the model defaults to generating a parrot closer to the prompted red color. The lower row demonstrates this effect across a broader range of colors. Pink succeeds due to shared properties with red (both are warm colors with similar hue angles), while semantically distant colors (olive, navy, yellow) fail to override red descriptors.
@@ 5 Conclusion
We introduced ColorWave, the first training-free method for precise color control in text-to-image diffusion models. By exploiting latent links between visual attributes and semantic cues in IP-Adapter, our approach enables direct RGB-level manipulation without fine-tuning. It rewires cross-attention to achieve accurate color attribution while preserving generation quality. Extensive evaluations show ColorWave outperforms training-free methods in color accuracy and rivals fine-tuned approaches like ColorPeel—offering faster, more efficient alternatives to prompt-based color control. This reveals new, untapped capabilities in existing diffusion models.
Acknowledgments. This work was supported by Grants PID2021-128178OB-I00, PID2022-143257NB-I00, and PID2024-162555OB-I00 funded by MCIN/AEI/10.13039/ 501100011033 and FEDER, by the Generalitat de Catalunya CERCA Program, by the grant Catedra ENIA UAB-Cruïlla (TSI-100929-2023-2) from the Ministry of Economic Affairs and Digital Transition of Spain. JVC also acknowledges the 2025 Leonardo Grant for Scientific Research and Cultural Creation from the BBVA Foundation. The BBVA Foundation accepts no responsibility for the opinions, statements and contents included in the project and/or the results thereof, which are entirely the responsibility of the authors. Kai Wang acknowledges the funding from Guangdong and Hong Kong Universities 1+1+1 Joint Research Collaboration Scheme and the start-up grant B01040000108 from CityU-DG. The authors thankfully acknowledges RES resources provided by BSC in MareNostrum5 to IM-2025-3-0025, IM-2025-3-0027.
@@ Appendix A Color binding regions as a function of steps
Figure 11 displays color binding regions based on $\delta_{\lambda}(\tilde{\mathbf{A}}_{o})$ every 10-th denoising step,
Figure 11: Color binding regions based on $\delta_{\lambda}(\tilde{\mathbf{A}}_{o})$. Depicted every 10-th denoising step.
Appendix B Multi-object
ColorWave successfully applies different colors to multiple objects within the same scene by leveraging separate attention maps for each target object token, as demonstrated in Figure 12. As seen, the method achieves a robust color control. Potential failure cases may arise when the object has faulty or misrepresented attention maps, therefore rendering the color binding difficult or unfaithful.
Figure 12: Multi-object color control examples. (first) “A green rose growing from a red vase on a table”, (second) “A woman wearing blue pants sitting on a white sofa”, (third) “A blue teddy bear in Times Square on top of a pink car”, (fourth) “An orange parrot with a purple hat perched on a tree”, (fifth) “A photo of a blue dog on a yellow surfboard surfing a wave wearing a orange lifevest”.
Appendix C Experimental details
Evaluation metric details For each generated image, we use the Segment-Anything model [kirillov2023segment] to automatically extract masks for the target objects, allowing precise measurement of color attributes only within relevant regions. We report these metrics under multiple conditions: considering all pixels within the object mask (100%), and also considering the top 10% and 50% of pixels closest to the target color, which helps account for natural variations in object appearance such as highlights and shadows.
Implementation details We implement ColorWave using a pretrained Stable Diffusion XL (SDXL) [podell2023sdxl] model as our base architecture. For the image conditioning component, we utilize the IP-Adapter framework [ye2023ip] with an encoder based on OpenCLIP-ViT-H-14 [Ilharco_Open_Clip_2021]. We only inject the color embedding into the first decoder layer to compute cross-attention maps, which shows better stylization performance as previous works proved [wang2024instantstyle, wang2024instantstyle_plus, agarwal2023image_matte].
@@ Appendix C Experimental details
Appendix D Color Leakage Metric
To quantify unintended color attribution in generated images, we introduce a color leakage metric that measures the percentage of pixels in non-target regions that exhibit colors similar to the specified target color. This metric is particularly important for evaluating whether color control methods successfully confine the desired color to the intended object without affecting other parts of the image.
The color leakage metric operates by first extracting the reference RGB color from the image filename, then converting both the reference color and the generated image to HSV color space. For each pixel in the image, we calculate the circular hue difference between the pixel’s hue and the reference hue, accounting for the wraparound nature of the hue channel at $180^{\circ}$. Pixels with hue differences within a threshold of 10 degrees are considered to match the target color.
Appendix E Mechanism of Semantic Attribute Binding
As depicted in Figure 13, the semantic attribute binding phenomenon arises from IP-Adapter’s decoupled cross-attention design, where shared query projection matrices create implicit alignment between visual and textual feature spaces. We quantify this binding by computing the inner product similarity $\langle\mathbf{K}_{\text{text}},\mathbf{K}_{\text{image}}\rangle$ between text prompt tokens and image feature tokens.
The heatmap reveals systematic correspondence patterns across color combinations, with diagonal peaks indicating that the model implicitly learns to associate linguistic color terms with their visual counterparts. The token-level analysis demonstrates that color words (e.g., ”pink”) exhibit peak similarity with corresponding visual color features, while non-color tokens maintain lower similarities. This mechanism generalizes across the continuous color space, with perceptually similar colors showing elevated cross-similarities that reflect the model’s understanding of color relationships.
Figure 13: The dual cross-attention architecture enables implicit alignment between text and image features through shared query projections. The similarity heatmap (bottom left) shows correspondence patterns between prompt colors and target colors, with diagonal peaks indicating successful color-semantic binding. The token-level similarity plot (bottom right) demonstrates how individual prompt tokens relate to image features, with color words showing peak correspondence to matching visual color information.
Appendix F Extra results of ColorWave
Extra results are showcased in Figures 14 to 21.